# SLS-BRD: A system-level approach to seeking generalised feedback Nash equilibria

### v2

@@ III-A System-level best-response mappings
Collectively, the mapping $BR_{\Phi}:\bm{\mathcal{C}}_{u}\rightrightarrows\bm{\mathcal{C}}_{u}$, defined by $BR_{\Phi}(\bm{\Phi_{u}})=BR^{1}_{\Phi}(\bm{\Phi_{u}}^{-1})\times\cdots\times BR^{N_{P}}_{\Phi}(\bm{\Phi_{u}}^{-N_{P}})$, is the joint best-response to a system-level strategy profile $\bm{\Phi_{u}}$. As before, the GNEs of $\mathcal{G}_{\infty}^{\Phi}$ are equivalent to the fixed-points of this map, $\Omega_{\mathcal{G}_{\infty}^{\Phi}}=\{\bm{\Phi_{u}}^{*}\in\bm{\mathcal{C}}_{u}:\bm{\Phi_{u}}^{*}\in BR_{\Phi}(\bm{\Phi_{u}}^{*})\}$. Considering how $\mathcal{G}_{\infty}^{\text{LQ}}$ induces $\mathcal{G}_{\infty}^{\Phi}$, a relationship can be established between the best-responses $BR$ and $BR_{\Phi}$ and, consequently, between their fixed-points, $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$ and $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$. In Section III-B, we formalise this relationship and propose a learning dynamics for GFNE seeking based on the system-level best-response mappings.
Remark 1.
Problem (31) is independent of the matrix $\Sigma_{w}^{1\over 2}$ due to its linear-quadratic structure and the fact that $\Sigma_{w}\in\mathbb{S}^{N_{x}}$. Hence, we remove it from $J^{p}$ ($\forall p$) to simplify notation.
The best-responses $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ are still intractable: i) They are defined by infinite-dimensional problems with no general solution and ii) that require full knowledge of the noise process $(w_{n})_{n\in\mathbb{N}}$ to formulate the constraints $U_{\Phi}^{p}$. In the following, we tackle both issues and provide a class of finite-dimensional robust optimisation programs that approximate Problem (31). We conclude the section by presenting a class of system level constraints which enforce a richer feedback information pattern.
@@
The programs in $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ can be made finite-dimensional by restricting the closed-loop maps to the set of finite-impulse responses (FIR),
$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}\in\ell_{2}[0,N]:\Phi_{x,n}\in\mathcal{C}_{x,n},\leavevmode\nobreak\ n{\in}[0,N),\leavevmode\nobreak\ \Phi_{x,N}=0\};$
$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}\in\ell_{2}[0,N):\Phi_{u,n}^{p}\in\mathcal{C}^{p}_{u,n},\leavevmode\nobreak\ n{\in}[0,N)\},$
given horizon $N\in[2,\infty)$. We enforce $\bm{\Phi_{x}}\in\bm{\mathcal{C}}_{x}$ and $\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}$ in Problem (31) by adding the terminal constraint $\Phi_{x,N}=0$, and restricting the spectral factors to satisfy $(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N-1}$. The constraint sets $\mathcal{C}_{x,n}\subseteq\mathbb{R}^{N_{x}\times N_{x}}$ and $\mathcal{C}_{u,n}^{p}\subseteq\mathbb{R}^{N_{u}\times N_{x}}$ ($\forall p$) have no restriction other than being compact convex sets. The resulting problems are of the form
$\underset{\bm{\Phi_{u}}^{p}=(\Phi_{u,n}^{p})_{n=1}^{N-1}}{\text{minimize}}\quad$ $\sum_{n=1}^{N}\Big{(}\|C^{p}\Phi_{x,n}\|_{F}^{2}+\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\|_{F}^{2}\Big{)}$ (32a)
$\underset{\forall n\in[1,N)}{\text{subject to}}\quad\leavevmode\nobreak\ \leavevmode\nobreak$ $\Phi_{x,{n+1}}=A\Phi_{x,n}+\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (32b)
@@
$\Phi_{x,n}\in\mathcal{C}_{x,n},\quad\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (32f)
$\Phi_{x,1}=I_{N_{x}},\quad\Phi_{x,N}=0,$ (32g)
which define finite-dimensional convex programs that can be solved numerically. Since $(\bm{\mathcal{C}}_{x},\bm{\mathcal{C}}_{u}^{p})$ are finite-dimensional, the operators $(\bm{F_{\Phi}},\bm{K})$ and the convolutions in Eq. (32e) can be represented by matrix multiplications: Consider any kernel $\bm{\Phi}\in\{\bm{\Phi_{x}}\}\cup\{\bm{\Phi}^{p},\bm{\Phi_{u}}^{p}\}_{p\in\mathcal{P}}$ and some signal $\bm{z}\in\ell_{\infty}$. Then,
$\bm{z}^{\text{out}}=\bm{\Phi}*\bm{z}\leavevmode\nobreak\ \leavevmode\nobreak\ \Longleftrightarrow\leavevmode\nobreak\ \leavevmode\nobreak\ z^{\text{out}}=M_{\Phi}z,$
given block-Toeplitz matrix $M_{\Phi}=[\Phi_{i-j}]_{i\in[1,2N+1],j\in[1,N+1]}$ and $(z,z^{\text{out}})$ being a vector representation of these signals, that is, $z=\texttt{col}(z_{0},\ldots,z_{N})$ and $z^{\text{out}}=\texttt{col}(z^{\text{out}}_{0},\ldots,z^{\text{out}}_{2N})$. Moreover, $z^{\text{out}}_{n}=(\Phi*z)_{n}=[M_{\Phi}]_{n}z$, where $[M_{\Phi}]_{n}$ is the $n$-th block-row of $M_{\Phi}$. Thus, Problem (32) can be reformulated into a program requiring only matrix algebra. Finally, we remark that policies ($\bm{K}^{p}$, $p\in\mathcal{P}$) are implemented either by using Corollary 4.1 or by directly computing kernels $\bm{\Phi}^{p}=([M_{\Phi_{u}^{p}}]_{n}\Phi_{x}^{-1})_{n\in\mathbb{N}}$.
Although realising Problem (31) into a tractable program, the Problem (32) is only feasible when the pair $(A,B^{p})$ is full-state controllable. This is a difficult requirement in multi-agent settings, as often $N_{u}^{p}\ll N_{x}$ for all $p\in\mathcal{P}$, leading to Problem (32) becoming overdetermined. Furthermore, enforcing FIR constraints is known to result in deadbeat policies: Control actions are excessively large in magnitude for small $N<\infty$. Alternatively, we restrict the system level responses to the sets
$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}{\in}\ell_{2}[0,N]:\Phi_{x,n}{\in}\mathcal{C}_{x,n},\leavevmode\nobreak\ n{\in}[0,N),\leavevmode\nobreak\ \|\Phi_{x,N}\|_{F}^{2}\leq\gamma\};$
$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}{\in}\ell_{2}[0,N):\Phi_{u,n}^{p}{\in}\mathcal{C}^{p}_{u,n},\leavevmode\nobreak\ n{\in}[0,N)\},$
with $\|\Phi_{x,N}\|_{F}^{2}=\sum_{i}\sigma_{i}(\Phi_{x,N})^{2}\leq\gamma$ for some factor $\gamma\in(0,1)$ and $\sigma_{i}(\cdot)$ denoting the $i$-th largest singular value of a matrix. These are reffered to as the set of approximately FIR for $N>0$. The $p$-th player best-response map thus takes the form
$\underset{\bm{\Phi_{u}}^{p}=(\Phi_{u,n}^{p})_{n=1}^{N-1}}{\text{minimize}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \sum_{n=1}^{N-1}\Big{(}\|C^{p}\Phi_{x,n}\|_{F}^{2}{+}\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\|_{F}^{2}\Big{)}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \hskip 121.00024pt+\|C^{p}\Phi_{x,N}\|_{F}^{2}$ (33a)
$\underset{\forall n\in[1,N)}{\text{subject to}}$ $\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (33b)
[…]
$\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,n}\in\mathcal{C}_{x,n},\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (33f)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \Phi_{x,1}=I_{N_{x}},\leavevmode\nobreak\ \leavevmode\nobreak\ \|\Phi_{x,N}\|_{F}^{2}\leq\gamma.$ (33g)
@@
The equilibrium gap associated with $\bm{\Phi_{u}}^{\varepsilon}\in\widehat{BR}_{\Phi}(\bm{\Phi_{u}}^{\varepsilon})$ is thus proportional to the parameter $\gamma$, assuming that the terminal constraint Eq. (33g) also holds for solutions to the original Problem (31). Hereafter, $U_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ ($\forall p$) are assumed to incorporate the pair ($\bm{\mathcal{C}}_{x}$, $\bm{\mathcal{C}}_{u}^{p}$) defining the set of (soft-constrained) FIR approximations defined above for a $N>0$. We write $U_{\Phi,N}^{p}$ whenever this needs to be made explicit (e.g., to distinguish this choice from a more general $\bm{\mathcal{C}}_{x}$ and $\bm{\mathcal{C}}_{u}^{p}$).
@@
From Assumption 3, the sets $\mathcal{X}$, $\mathcal{U}^{p}$, and $\mathcal{U}_{\mathcal{G}}$, can be expressed by linear inequalities,
$\mathcal{X}$ =$=\{x_{t}\in\mathbb{R}^{N_{x}}:G_{x}x_{t}\leq g_{x}\};$
$\mathcal{U}^{p}$ =$=\{u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:G_{u}^{p}u^{p}_{t}\leq g_{u}^{p}\};$
$\mathcal{U}_{\mathcal{G}}$ =$=\{u_{t}\in\textstyle\prod_{\tilde{p}\in\mathcal{P}}\mathbb{R}^{N_{u}^{\tilde{p}}}:\textstyle\sum_{\tilde{p}\in\mathcal{P}}G_{\mathcal{G}}^{\tilde{p}}u^{\tilde{p}}_{t}\leq g_{\mathcal{G}}\},$
@@
given some matrices $G_{x}\in\mathbb{R}^{N_{\mathcal{X}}\times N_{x}}$, $G_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}^{p}}\times N_{u}^{p}}$ and $G_{\mathcal{G}}^{\tilde{p}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}\times N_{u}^{\tilde{p}}}$ ($\forall\tilde{p}\in\mathcal{P}$), and vectors $g_{x}\in\mathbb{R}^{N_{\mathcal{X}}}_{\geq 0}$, $g_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}}^{p}}_{\geq 0}$, and $g_{\mathcal{G}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}}_{\geq 0}$. The map $U^{p}(u^{-p})$ is then equivalent to the actions $\bm{u}^{p}=\bm{\Phi_{u}}^{p}*\bm{w}$ whose associated response $\bm{\Phi_{u}}^{p}$ satisfy
$[G_{x}]_{i}(\Phi_{x}*w)_{n}$ $\leq[g_{x}]_{i},\quad i=1,\ldots,N_{\mathcal{X}};$ (35)
$[G_{u}^{p}]_{j}(\Phi_{u}^{p}*w)_{n}$ $\leq[g_{u}^{p}]_{j},\quad j=1,\ldots,N_{\mathcal{U}^{p}};$ (36)
$\textstyle\sum_{\tilde{p}\in\mathcal{P}}[G_{\mathcal{G}}^{\tilde{p}}]_{l}(\Phi_{u}^{\tilde{p}}*w)_{n}$ $\leq[g_{\mathcal{G}}]_{l},\quad l=1,\ldots,N_{\mathcal{U}_{\mathcal{G}}},$ (37)
@@
with $([G_{x}]_{i},[g_{x}]_{i})$, $([G_{u}^{p}]_{j},[g_{u}^{p}]_{j})$, and $([G_{\mathcal{G}}^{\tilde{p}}]_{l},[g_{\mathcal{G}}]_{l})$ denoting, respectively, the $i$-th, $j$-th, and $l$-th rows of the corresponding matrices and vectors. In this work, players are assumed to synthesise policies that satisfy these constraints for any $\bm{w}\in\bm{\mathcal{W}}$. At the expense of conservativeness, we cast Problem (33) as a robust optimization problem by considering the worst-case realisation of the noise. Specifically, we reformulate Eq. (35),
$\sum_{n^{\prime}=0}^{N}\sup_{w\in\mathcal{W}}[G_{x}]_{i}\Phi_{x,n^{\prime}}w\leq[g_{x}]_{i},\quad i=1,\ldots,N_{\mathcal{X}}$ (38)
and similarly for Eqs. (36)–(37), then exploit our knowledge of $\bm{\mathcal{W}}$ to obtain an analytical solution for the supremum.
A common instance of $\mathcal{G}_{\infty}^{\text{LQ}}$ considers the problem in which $(w_{n})_{n\in\mathbb{N}}$ is uniformly bounded. In these problems, the inequality Eq. (38) can be expressed in terms of a dual norm on the appropriate vector space. We highlight two important cases.
  • $\mathcal{W}$ is an ellipsoid centred at zero. Formally, the set is
    $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|Pw_{t}\|_{2}\leq 1\},$
    given a matrix $P\in\mathbb{S}_{++}^{N_{x}}$. This corresponds to noise processes with uniformly bounded energy, as encoded by $\|\cdot\|_{2}$. For such cases, Eqs. (35)–(37) can be enforced by the second-order conic (SOC) constraints
    $\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{x}]_{i},\quad(\forall i);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{u}^{p}]_{j},\quad(\forall j);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}}P^{-1})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{2}$ $\leq[g_{\mathcal{G}}]_{l},\quad(\forall l).$
  • $\mathcal{W}$ is a polyhedron, symmetric around zero. Formally,
    $\mathcal{W}=\{w_{t}\in\mathbb{R}^{N_{x}}:\|Pw_{t}\|_{\infty}\leq 1\},$
    given a full-rank matrix $P\in\mathbb{R}^{N_{\mathcal{W}}\times N_{x}}$. This corresponds to noise processes with uniformly bounded intensity, as encoded by $\|\cdot\|_{\infty}$. For such cases, Eqs. (35)–(37) can be enforced by the first-order conic constraints
    $\textstyle\sum_{n=0}^{N}\big{\|}([G_{x}]_{i}\Phi_{x,n}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{x}]_{i},\quad(\forall i);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}([G_{u}^{p}]_{j}\Phi_{u,n}^{p}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{u}^{p}]_{j},\quad(\forall j);$
    $\textstyle\sum_{n=0}^{N{-}1}\big{\|}\textstyle\sum_{\tilde{p}\in\mathcal{P}}([G_{\mathcal{G}}^{\tilde{p}}]_{l}\Phi_{u,n}^{\tilde{p}}P^{\dagger})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{1}$ $\leq[g_{\mathcal{G}}]_{l},\quad(\forall l).$
@@
If $\mathcal{X}=\mathbb{R}^{N_{x}}$, $\mathcal{U}^{p}=\mathbb{R}^{N_{u}^{p}}$, and $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$, these constraints are trivially satisfied for any $w\in\bm{\mathcal{W}}$, and thus can be removed from Problem (31). Conversely, if $\mathcal{W}=\mathbb{R}^{N_{x}}$ when either $\mathcal{X}$, $\mathcal{U}^{p}$, or $\mathcal{U}_{\mathcal{G}}$ is bounded, then no stabilising policy can enforce those constraints for all possible $\bm{w}$.

### v3

@@ 3.1 System-level best-response mappings
We refer to the Supplementary Material for a detailed derivation of Problem (29) from Problem (17). Collectively, the mapping $BR_{\Phi}(\bm{\Phi_{u}})=BR^{1}_{\Phi}(\bm{\Phi_{u}}^{-1})\times\cdots\times BR^{N_{P}}_{\Phi}(\bm{\Phi_{u}}^{-N_{P}})$, is the joint best-response to a system-level strategy profile $\bm{\Phi_{u}}$. The GNEs of $\mathcal{G}_{\infty}^{\Phi}$ are equivalent to the fixed-points of this map, $\Omega_{\mathcal{G}_{\infty}^{\Phi}}=\{\bm{\Phi_{u}}^{\star}\in\bm{\mathcal{C}}_{u}:\bm{\Phi_{u}}^{\star}\in BR_{\Phi}(\bm{\Phi_{u}}^{\star})\}$. Considering how $\mathcal{G}_{\infty}^{\text{LQ}}$ induces $\mathcal{G}_{\infty}^{\Phi}$, a relationship can be established between the best-responses $BR$ and $BR_{\Phi}$ and, consequently, between their fixed-points, $\Omega_{\mathcal{G}_{\infty}^{\text{LQ}}}^{\bm{K}}$ and $\Omega_{\mathcal{G}_{\infty}^{\Phi}}$. In Section 3.2, we formalise this relationship and propose a learning dynamics for GFNE seeking based on the system-level best-response mappings.
The best-responses $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ are still intractable: i) They are defined by infinite-dimensional problems with no general solution and ii) that require full knowledge of the noise process $(w_{n})_{n\in\mathbb{N}}$ to formulate the constraints $U_{\Phi}^{p}$. In the following, we tackle both issues and provide a class of finite-dimensional robust optimisation problems that approximate Problem (29). We conclude the section by presenting a class of system level constraints which enforce a richer feedback information pattern.
@@
The programs in $\{BR_{\Phi}^{p}\}_{p\in\mathcal{P}}$ can be made finite-dimensional by restricting the system level responses to the set of finite-impulse responses (FIR),
$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}\in\ell_{2}[0,N]:\Phi_{x,n}\in\mathcal{C}_{x,n},~{}n{\in}[0,N),~{}\Phi_{x,N}=0\};$
$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}\in\ell_{2}[0,N):\Phi_{u,n}^{p}\in\mathcal{C}^{p}_{u,n},~{}n{\in}[0,N)\},$
given horizon $N\in(1,\infty)$. We enforce $\bm{\Phi_{x}}\in\bm{\mathcal{C}}_{x}$ and $\bm{\Phi_{u}}^{p}\in\bm{\mathcal{C}}_{u}^{p}$ in Problem (29) by adding a terminal constraint $\Phi_{x,N}=0$ then letting $\bm{\Phi_{x}}=(\Phi_{x,n}\in\mathcal{C}_{x,n})_{n=1}^{N}$ and $\bm{\Phi_{u}}=(\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p})_{n=1}^{N-1}$. The sets $\mathcal{C}_{x,n}\subseteq\mathbb{R}^{N_{x}\times N_{x}}$ and $\mathcal{C}_{u,n}^{p}\subseteq\mathbb{R}^{N_{u}^{p}\times N_{x}}$ are required only to be compact convex sets. Under this condition, Problem (29) is finite-dimensional with $(N{-}1)N_{x}N_{u}^{p}$ decision variables (the entries of $\Phi_{u,1}^{p},\ldots,\Phi_{u,N{-}1}^{p}\in\mathbb{R}^{N_{u}^{p}\times N_{x}}$) and thus can be solved numerically. We remark that the policies $\bm{K}^{p}=\bm{\Phi_{u}}^{p}\bm{\Phi_{x}}^{-1}$ ($p\in\mathcal{P}$) are still solutions to the infinite-horizon Problem (17) regardless of the closed-loop maps $\{\bm{\Phi_{x}},\bm{\Phi_{u}}^{p}\}$ being FIR.
Although realising Problem (29) into a tractable program, the constraint $\Phi_{x,N}=0$ is only feasible when the pair $(A,B^{p})$ is full-state controllable. This is a difficult requirement in multi-agent settings, as often $N_{u}^{p}\ll N_{x}$ for all $p\in\mathcal{P}$, leading to overdetermined problems. Furthermore, enforcing FIR constraints is known to result in deadbeat policies: Control actions are excessively large in magnitude for small $N<\infty$. Alternatively, we restrict the system level responses to the sets
$\bm{\mathcal{C}}_{x}$ =$=\{\bm{\Phi_{x}}{\in}\ell_{2}[0,N]:\Phi_{x,n}{\in}\mathcal{C}_{x,n},~{}n{\in}[0,N),~{}\|\Phi_{x,N}\|_{F}^{2}\leq\gamma\};$
$\bm{\mathcal{C}}_{u}^{p}$ =$=\{\bm{\Phi_{u}}^{p}{\in}\ell_{2}[0,N):\Phi_{u,n}^{p}{\in}\mathcal{C}^{p}_{u,n},~{}n{\in}[0,N)\},$
with $\|\Phi_{x,N}\|_{F}^{2}=\sum_{i}\sigma_{i}(\Phi_{x,N})^{2}\leq\gamma$ for some factor $\gamma\in(0,1)$ and $\sigma_{i}(\cdot)$ denoting the $i$-th largest singular value of a matrix. These are denoted as the set of (soft) FIR for $N>0$. The $p$-th player’s best-response map thus corresponds to the problem
$\underset{\bm{\Phi_{u}}^{p}}{\text{minimize}}$ $~{}~{}\sum_{n=1}^{N-1}\Big{(}\|C^{p}\Phi_{x,n}\Sigma_{w}^{1\over 2}\|_{F}^{2}{+}\|\textstyle\sum_{\tilde{p}\in\mathcal{P}}D^{p\tilde{p}}\Phi_{u,n}^{\tilde{p}}\Sigma_{w}^{1\over 2}\|_{F}^{2}\Big{)}$
$~{}~{}\hskip 124.00012pt+\|C^{p}\Phi_{x,N}\Sigma_{w}^{1\over 2}\|_{F}^{2}$ (30a)
$\underset{\forall n\in[1,N)}{\text{subject to}}$ $~{}~{}\Phi_{x,{n+1}}=A\Phi_{x,n}{+}\textstyle\sum_{\tilde{p}\in\mathcal{P}}B^{\tilde{p}}\Phi^{\tilde{p}}_{u,n},$ (30b)
@@
$~{}~{}\Phi_{x,n}\in\mathcal{C}_{x,n},~{}~{}\Phi_{u,n}^{p}\in\mathcal{C}_{u,n}^{p},$ (30f)
$~{}~{}\Phi_{x,1}=I_{N_{x}},~{}~{}\|\Phi_{x,N}\|_{F}^{2}\leq\gamma.$ (30g)
@@
The complexity of Problem (30) is agnostic to the number of players: The effect of other players’ strategies can always be condensed as affine terms (e.g., $z^{-p}_{n}=\sum_{\tilde{p}\in\mathcal{P}\backslash\{p\}}B^{\tilde{p}}\Phi_{u,n}^{\tilde{p}}$). Conversely, it scales quickly with the FIR horizon $N$ and model dimensions $N_{x}$ and $N_{u}^{p}$. We remark, however, that this problem is still convex and can be solved efficiently by exploiting its structure using techniques from real-time optimisation of linear-quadratic control problems [42]. Moreover, if the constraints Eq. (30e)–(30g) are column-separable (see [35]), then Problem (30) can be decomposed into smaller subproblems to be solved in parallel, substantially reducing its computational costs.
@@
The equilibrium gap associated with $\bm{\Phi_{u}}^{\varepsilon}\in\widehat{BR}_{\Phi}(\bm{\Phi_{u}}^{\varepsilon})$ thus depends on the choice of parameter $\gamma$, assuming that the terminal constraint Eq. (30g) also holds for solutions to the original Problem (29). Hereafter, $U_{\Phi}^{p}:\bm{\mathcal{C}}_{u}^{-p}\rightrightarrows\bm{\mathcal{C}}_{u}^{p}$ ($\forall p$) are assumed to incorporate the sets ($\bm{\mathcal{C}}_{x}$, $\bm{\mathcal{C}}_{u}^{p}$) of (soft-constrained) FIR approximations as defined above for a $N>1$.
@@
From Assumption 3, the sets $\mathcal{X}$, $\mathcal{U}^{p}$ ($p\in\mathcal{P}$), and $\mathcal{U}_{\mathcal{G}}$, can be expressed by linear inequalities,
$\mathcal{X}$ =$=\{x_{t}\in\mathbb{R}^{N_{x}}:G_{x}x_{t}\preceq\mathbf{1}_{N_{\mathcal{X}}}\};$
$\mathcal{U}^{p}$ =$=\{u^{p}_{t}\in\mathbb{R}^{N_{u}^{p}}:G_{u}^{p}u^{p}_{t}\preceq\mathbf{1}_{N_{\mathcal{U}}^{p}}\};$
$\mathcal{U}_{\mathcal{G}}$ =$=\{u_{t}\in\textstyle\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}:G_{\mathcal{G}}u_{t}\preceq\mathbf{1}_{N_{\mathcal{U}_{\mathcal{G}}}}\},$
@@
given some matrices $G_{x}\in\mathbb{R}^{N_{\mathcal{X}}\times N_{x}}$, $G_{u}^{p}\in\mathbb{R}^{N_{\mathcal{U}^{p}}\times N_{u}^{p}}$, and $G_{\mathcal{G}}\in\mathbb{R}^{N_{\mathcal{U}_{\mathcal{G}}}\times N_{u}}$. The map $U^{p}(u^{-p})$ is then equivalent to the actions $\bm{u}^{p}=\bm{\Phi_{u}}^{p}*\bm{w}$ whose associated response $\bm{\Phi_{u}}^{p}$ satisfy
$[G_{x}]_{i}(\Phi_{x}*w)_{n}$ $\leq 1,\quad i=1,\ldots,N_{\mathcal{X}};$ (32)
$[G_{u}^{p}]_{j}(\Phi_{u}^{p}*w)_{n}$ $\leq 1,\quad j=1,\ldots,N_{\mathcal{U}^{p}};$ (33)
$[G_{\mathcal{G}}]_{l}(\Phi_{u}*w)_{n}$ $\leq 1,\quad l=1,\ldots,N_{\mathcal{U}_{\mathcal{G}}},$ (34)
@@
with $([G_{x}]_{i},[G_{u}^{p}]_{j},[G_{\mathcal{G}}]_{l})$ being the $i$-th, $j$-th, and $l$-th rows of the corresponding matrices. In this work, players are assumed to synthesise policies satisfying these constraints for any $\bm{w}\in\bm{\mathcal{W}}$. Equivalently, we cast Problem (30) as a robust optimization problem by considering the worst-case realisation of the noise. Specifically, we reformulate the constraint Eq. (32) as
$\sup_{\bar{w}\in\mathcal{W}^{N}}\left\{\textstyle\sum_{n^{\prime}=0}^{N}[G_{x}]_{i}\Phi_{x,n^{\prime}}\bar{w}_{n^{\prime}}\right\}\leq 1,\quad i=1,\ldots,N_{\mathcal{X}},$
and similarly for Eqs. (33)–(34). Since $(\bm{\Phi}_{x},\bm{\Phi}_{u}^{p})$ are FIR maps, it suffices to consider noise sequences of length $N$, $\bar{w}\in\mathcal{W}^{N}=\prod_{n^{\prime}=1}^{N}\mathcal{W}$, then exploit our knowledge of $\mathcal{W}$ to analytically solve this supremum. A common instance of $\mathcal{G}_{\infty}^{\text{LQ}}$ consists on the problems in which $(w_{n})_{n\in\mathbb{N}}$ is known to satisfy
$w_{n}\in\mathcal{W}=\{P\zeta\in\mathbb{R}^{N_{x}}:\|\zeta\|_{q}\leq 1\},\quad\forall n\in\mathbb{N},$
given a full-column-rank matrix $P\in\mathbb{R}^{N_{x}\times N_{\zeta}}$ and the $\ell_{q}$-norm $\|\cdot\|_{q}$. Using standard results from linear algebra [43], the robust counterpart of constraints Eqs. (32)–(34) are the constraints
$N^{1/q}\big{\|}(I_{N}\otimes P^{{\mkern-1.5mu\mathsf{T}}})([G_{x}]_{i}\bar{\Phi}_{x})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{q}^{*}$ $\leq 1,\quad(\forall i);$ (35)
$(N{-}1)^{1/q}\big{\|}(I_{N{-}1}\otimes P^{{\mkern-1.5mu\mathsf{T}}})([G_{u}^{p}]_{j}\bar{\Phi}_{u}^{p})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{q}^{*}$ $\leq 1,\quad(\forall j);$ (36)
$(N{-}1)^{1/q}\big{\|}(I_{N{-}1}\otimes P^{{\mkern-1.5mu\mathsf{T}}})([G_{\mathcal{G}}]_{l}\bar{\Phi}_{u})^{{\mkern-1.5mu\mathsf{T}}}\big{\|}_{q}^{*}$ $\leq 1,\quad(\forall l),$ (37)
in which $\bar{\Phi}_{x}=[\Phi_{x,1}~{}\cdots~{}\Phi_{x,N}]$, $\bar{\Phi}_{u}^{p}=[\Phi_{u,1}^{p}~{}\cdots~{}\Phi_{u,N-1}^{p}]$, and $\bar{\Phi}_{u}=[\Phi_{u,1}~{}\cdots~{}\Phi_{u,N-1}]$. We refer to the Supplementary Material for a detailed derivation. The Problem (30) is thus rendered robust to the uncertainty in $\bm{w}\in\bm{\mathcal{W}}$ by incorporating the worst-case constraints Eqs. (35)–(37) in place of Eq. (30e).
@@
If $\mathcal{X}=\mathbb{R}^{N_{x}}$, $\mathcal{U}^{p}=\mathbb{R}^{N_{u}^{p}}$, or $\mathcal{U}_{\mathcal{G}}=\prod_{p\in\mathcal{P}}\mathbb{R}^{N_{u}^{p}}$, then the corresponding constraints are trivially satisfied for any $w\in\bm{\mathcal{W}}$ and can be removed from Problem (29). Conversely, if $\mathcal{W}=\mathbb{R}^{N_{x}}$ when either $\mathcal{X}$, $\mathcal{U}^{p}$, or $\mathcal{U}_{\mathcal{G}}$ is bounded, then no stabilising policy can enforce those constraints for all $\bm{w}$.
Remark 2.
The constraints Eq. (35)–(37) are equivalent to those with $(-[G_{x}]_{i},-[G_{u}^{p}]_{j},-[G_{\mathcal{G}}]_{l})$. As such, matrices in the form $\widetilde{G}=\texttt{col}(G,-G)$ (i.e., enforcing $-\bm{1}_{N_{z}}\preceq Gz\preceq\bm{1}_{N_{z}}$) can be replaced by $G$ ($Gz\preceq\bm{1}_{N_{z}}$) and yield the same results.